OpenStudy (starryknight):

Find the sum of the first 26 terms of the following sequence: 12, 18, 24 . . .

OpenStudy (starryknight):

i got 2210 but idont know if its right or wrong

hartnn (hartnn):

what did you get for common difference ?

OpenStudy (starryknight):

12

OpenStudy (starryknight):

actually -6

hartnn (hartnn):

common difference = d = difference between consecutive terms d= 18-12 = 24-18 = ... ?

hartnn (hartnn):

not even -6

OpenStudy (starryknight):

sorry got mixed up 6

hartnn (hartnn):

now use the sum formula. have it ?

hartnn (hartnn):

n =26 a= 12 d=6

OpenStudy (starryknight):

s26=26/2(2*2(26-1)6) 26/2(2*2(25)6) 26/2(2*12+(150) 26/2(24+(150) 26/2(174) 13(1740 2262

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (starryknight):

thanx again :)

hartnn (hartnn):

welcome ^_^